# Normalized topological indices discriminate between architectures of branched macromolecules

### v1

@@ Abstract
Branching architecture characterizes numerous systems, ranging from synthetic (hyper)branched polymers and biomolecules such as lignin, amylopectin, and nucleic acids to tracheal and neuronal networks. Its ubiquity reflects the many favourable properties that arise because of it. For instance, branched macromolecules are spatially compact and have a high surface functionality, which impacts their phase characteristics and self-assembly behaviour, among others. The relationship between branching and physical properties has been studied by mapping macromolecules to mathematical trees whose architecture can be characterized using topological indices. These indices, however, do not allow for a comparison of macromolecules that map to trees of different size, be it due to different mapping procedures or differences in their molecular weight. To alleviate this, we introduce a novel normalization of topological indices using estimates of their probability density functions. We determine two optimal normalized topological indices and construct a phase space that enables a robust discrimination between different architectures of branched macromolecules. We demonstrate the necessity of such a phase space on two practical applications, one being ribonucleic acid (RNA) molecules with various branching topologies and the other different methods of coarse-graining branched macromolecules. Our approach can be applied to any type of branched molecules and extended as needed to other topological indices, making it useful across a wide range of fields where branched molecules play an important role, including polymer physics, green chemistry, bioengineering, biotechnology, and medicine.
@@ Introduction
In this work, we introduce a novel normalization of topological indices based on estimating their probability density functions. We determine two optimal normalized topological indices and construct a phase space that enables robust discrimination between the topologies of molecules of different types and sizes. The necessity of our approach is illustrated with two practical examples. First, we use the phase space to distinguish between topologies of RNA secondary structures with different degrees of branching and to show how different tree representations of RNA influence the resulting topology. Second, we compare two methods of coarse-graining the structures of branched macromolecules and use the phase space of normalized topological indices to highlight the differences between the two methods. Taken together, our work establishes a phase space based on two normalized topological indices that allows for a well-defined way of discriminating between branching architectures of different macromolecules as well as of methods of coarse-graining them.
@@
Results
Tree representation of branched macromolecules
Branching is characteristic of macromolecules across different systems and scales (Fig. 1). To compare their topologies, one can first map their structure onto a mathematical tree $T$ with $N$ edges, with $V(T)$ denoting the set of its nodes and $E(T)$ denoting the set of its edges. This mapping is generally not unique and depends on the macromolecule [44]. For instance, a molecular (chemical) graph has atoms as nodes and edges that represent chemical bonds (Fig. 1a). Due to the nature of carbon bonding, chemical graphs have a maximum branching degree of $4$ regardless of their size. For polymer structures, a typical tree representation is one where a node is a collection of atoms at a polymer branch point and an edge is a linear structure formed by a chain of bonds between two such branch points (Fig. 1b). A more complex case, RNA is a linear polymer whose constituent nucleotides interact and base-pair with one another and give rise to a complex and highly branched secondary structure. This structure is often mapped to a tree by associating the base-paired (double-stranded) regions with edges and unpaired (single-stranded) regions with nodes (Fig. 1c). As we will see later on, such a coarse-grained fashion of mapping macromolecules to trees can lead to RNA trees with significantly larger branching degrees compared to chemical graphs.
Constructing the phase space of topological indices
Once a branched macromolecule is mapped to a mathematical tree, its architecture can be studied using methods from graph theory. This includes a wide variety of topological indices designed to describe the topology of trees and, more generally, graphs, with applications ranging from the description of chemical properties of molecules to connectivity of networks [45, 46, 47]. Broadly speaking, we can divide topological indices into path-, degree-, centrality-, spectrum-, and information-based indices depending on the main graph property used in their definition. For simplicity, we focus here on $18$ commonly used topological indices; their definitions are given in Supporting Information (SI). Since most topological indices were designed to be applied to graphs in general, which include cycles, they tend to be highly correlated when used on trees (Fig. S1 in SI). Individual indices are also highly degenerate and do not discriminate well between different tree topologies [51]. Therefore, our first aim is to select a pair of topological indices which are correlated as little as possible so that we can use their phase space for comparison of tree topologies.
We find that out of the $18$ topological indices we examined, one of the least correlated pairs (Fig. S1) is formed by second-order network coherence $H_{2}$ and atom bond connectivity index $\mathrm{ABC}$. The second-order network coherence is spectrum-based and defined as
@@ Results
The phase space formed by the pair of indices $(\mathrm{ABC},H_{2})$ permits us to discriminate between tree topologies at fixed tree size $N$ (Fig. 2a) significantly better than if we would have used either of the two indices alone (Fig. S6 in SI). We can furthermore trace the boundary of this phase space with special cases of tree topologies, ranging from linear tree and dumbbell trees to star tree, star-like trees, and spider tree (Fig. 2b), which provides us with a basic understanding of topologies found in different regions of the phase space.
Figure 2: Phase space constructed from topological indices ABC and $H_{2}$. (a) $(\mathrm{ABC},H_{2})$ space of $10^{4}$ random trees with $N=16$ edges. Distributions at the axes’ boundaries show the probability distribution functions of individual indices. (b) Different tree topologies that trace the boundary of the phase space, with large symbols in panels (a) and (d) corresponding to the shown tree structures. The trees at the boundary range from a linear tree, dumbbell trees, star tree, star-like trees, spider tree, and back to a linear tree; their construction is shown in Fig. S5 in SI. (c) $(\mathrm{ABC},H_{2})$ space for trees of four different sizes, where $10^{4}$ random trees are sampled for each $N$. (d) $(\mathrm{ABC}^{\prime},H_{2}^{\prime})$ space of trivially normalized indices [Eq. (3)] showing trees of four different sizes with $10^{4}$ random trees for each size (cf. panel (c)). (e) $(\mathrm{ABC}^{*},H_{2}^{*})$ space of indices normalized through their cumulative distribution functions [Eq. (5)] for $10^{4}$ random trees with $N=16$. Distributions at the axes’ boundaries show the probability distribution functions of individual indices. (f) $(\mathrm{ABC}^{*},H_{2}^{*})$ space of normalized indices for $N=30$, $50$, $100$, and $200$ with $10^{4}$ random trees for each tree size. (g) $(\mathrm{ABC}^{*},H_{2}^{*})$ space of all non-isomorphic trees with $N=16$, showing separately chemical and caterpillar trees as well as their union.
Comparing topologies of trees of different sizes
@@
Such normalization limits the index to the range $\mathrm{TI}^{\prime}\in[0,1]$, where $\mathrm{TI}^{\prime}(\mathrm{linear})=0$ and $\mathrm{TI}^{\prime}(\mathrm{star})=1$. However, this turns out to be too simple for the purpose of comparison, as it progressively compresses the phase space as $N$ increases (Fig. 2d) and makes discrimination between different branching architectures difficult. The compression of the phase space is due to the mean values of a topological index and its values for linear and star trees all scaling differently with tree size (Table S1 and Fig. S2); this effect is much more pronounced in the case of $H_{2}$ compared to $\mathrm{ABC}$.
@@
To determine the $\operatorname{cdf}(\mathrm{TI})$ of a given topological index, we need to first obtain its probability density function, $\operatorname{pdf}(\mathrm{TI})$. This is not a trivial endeavour since probability density functions for topological indices are not known. We therefore tested different continuous parametric distributions and chose the best-fitting one for each topological index (see Methods and Fig. S3 in SI). The $\operatorname{pdf}(\mathrm{ABC})$ is best described by a log-normal distribution [Eq. (6)], while Moyal distribution [Eq. (7)] is the best choice for the $\operatorname{pdf}(H_{2})$. The $N$-dependence of the fit parameters for both distributions is well-described by a simple power law and allows us to use the fitted $\operatorname{pdf}(\mathrm{TI})$ to immediately determine the cumulative distribution function $\operatorname{cdf}(\mathrm{TI})$ for an arbitrary tree size $N$.
The tree topologies assuming the extreme values of $\operatorname{cdf}(\mathrm{TI}(T))$ correspond to the trees that assume the extreme values of the original index $\mathrm{TI}$. If we want to define a normalized index $\mathrm{TI}^{*}$ so that its values for the linear and the star tree are fixed to be $\mathrm{TI}^{*}(\mathrm{linear})=0$ and $\mathrm{TI}^{*}(\mathrm{star})=1$, we have to normalize the cumulative distribution function accordingly. We thus define a normalized topological index $\mathrm{TI}^{*}$ as:
@@
Properties of $(\mathrm{ABC}^{*},H_{2}^{*})$ space
We can now use the normalization via cumulative distribution functions to create a space of normalized indices $(\mathrm{ABC}^{*},H_{2}^{*})$ which leads to a significantly more homogeneous phase space and enables us to compare topologies of trees of arbitrary size (panels (e) and (f) of Fig. 2). The $(\mathrm{ABC}^{*},H_{2}^{*})$ space also allows us to interpret the relationships between different tree topologies at its boundary (Fig. 2e) as well as inside it. For instance, chemical trees, which have a maximum branching degree of $4$, and caterpillar trees, whose nodes all lie within distance $1$ of a central path, occupy distinct parts of the phase space (Fig. 2g). The part of the phase space that chemical and caterpillar trees occupy furthermore shrinks with increasing tree size (panels (a) and (b) of Fig. S8 in SI) since their overall proportion compared to all trees decreases with increasing $N$. More generally, the position and proportion of different types of trees in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space is broadly determined by their maximum node degree and number of degree $2$ nodes. Specifically, the maximum node degree of a tree in the phase space increases as we go from a linear tree at $(0,0)$ to a star tree at $(1,1)$ (Fig. S8c); at the same time, the number of degree $2$ nodes diminishes (Fig. S8d).
We can also use the pair of normalized indices $(\mathrm{ABC}^{*},H_{2}^{*})$ to examine how different transformations of trees reposition their topology in the phase space. If we either move or remove a random leaf node, this repositions the new tree topology along the $\mathrm{ABC}$ axis but leaves its position along the $H_{2}$ axis fairly unchanged (panels (e) and (f) of Fig. S8). Transformations that affect larger parts of the tree topology—such as the Prüfer shuffle, which retains the node degree distribution of a tree but rearranges the connections between nodes, and random subtree swap, where an edge is randomly deleted and a new one randomly added—change the position of tree topology in the phase space more drastically (panels (g) and (h) of Fig. S8).
@@
Branching topologies of different tree representations of RNA structure
Base-pairing between the four constituent nucleotides of RNA leads to a branched secondary structure [21, 22] with numerous high-degree branching nodes ($d(i)>4$[20], making RNA unique among known branched macromolecules. This often carries functional significance: For instance, highly-branched RNA genomes of some viruses are compact and consequently more easily encapsidated during virion formation than less branched ones [56, 57]. Graphs present a natural way to coarse-grain RNA structure and analyze its branching topology [58]. However, the choice of the mapping of the base-pair pattern onto a mathematical tree is not straightforward and several options exist, designed with different purposes in mind [59, 60]. Perhaps the simplest way to map RNA secondary structure to a tree is by mapping double-stranded regions (base pairs) to edges, while single-stranded regions (unpaired nucleotides) are mapped to nodes connecting the edges [21, 22]. An alternative way, used in the RNA-as-graph (RAG) framework [58], omits from nodes those unpaired regions which consist of a single unpaired nucleotide or a single non-complementary pair of nucleotides, and does not consider lonely base pairs as edges. (For details on the two approaches, see Methods.)
@@
Secondary structures of the three RNA sequences are sampled from their thermal ensembles (Methods) and then mapped to trees using either the simple or the RAG method. While the RAG method (Fig. 3d) results in trees on average two thirds the size of those obtained using the simple method (Fig. 3a), the $(\mathrm{ABC}^{*},H_{2}^{*})$ space allows us to compare the RNA topologies resulting from the two mapping methods (panels (b) and (e) of Fig. 3). For both methods we observe that the three RNA sequences are separated in the phase space along the $H_{2}^{*}$ axis. The compact BlShV genome occupies large $H_{2}^{*}$ values, the non-compact BVQ genome occupies small $H_{2}^{*}$ values, and the random RNA sequence is located in-between the two. The $\mathrm{ABC}^{*}$ index, on the other hand, shows the differences between the two mapping methods. For the simple method, most trees acquire values of $\mathrm{ABC}^{*}$ close to zero. For the RAG method, however, the $\mathrm{ABC}^{*}$ values of both RNA genomes are distributed along the entire range of values; $\mathrm{ABC}^{*}$ values of the random RNA, while non-zero, remain small and in this way differ from both biological RNAs. These differences can be traced to the proportion of degree $2$ nodes produced by the two mapping methods, as the simple method (Fig. 3c) results in approximately twice as many degree $2$ nodes than the RAG method (Fig. 3f). As we have already observed previously, the amount of degree $2$ nodes is an important factor that determines the position of a tree topology in the phase space (Fig. S8d).
@@
Coarse-graining the topologies of random and scale-free trees

### v2

@@ Abstract
Branching architecture characterizes numerous systems, ranging from synthetic (hyper)branched polymers and biomolecules such as lignin, amylopectin, and nucleic acids to tracheal and neuronal networks. Its ubiquity reflects the many favourable properties that arise because of it. For instance, branched macromolecules are spatially compact and have a high surface functionality, which impacts their phase characteristics and self-assembly behaviour, among others. The relationship between branching and physical properties has been studied by mapping macromolecules to mathematical trees whose architecture can be characterized using topological indices. These indices, however, do not allow for a comparison of macromolecules that map to trees of different size, be it due to different mapping procedures or differences in their molecular weight. To alleviate this, we introduce a novel normalization of topological indices using estimates of their probability density functions. We construct a phase space using two normalized topological indices, which enables a robust discrimination between different architectures of branched macromolecules. We demonstrate the necessity of such a phase space on two practical applications, one being ribonucleic acid (RNA) molecules with various branching topologies and the other different methods of coarse-graining branched macromolecules. Our approach can be applied to any type of branched molecules and extended as needed to other topological indices, making it useful across a wide range of fields where branched molecules play an important role, including polymer physics, green chemistry, bioengineering, biotechnology, and medicine.
@@ I Introduction
In this work, we introduce a novel normalization of topological indices based on estimating their probability density functions. We choose two normalized topological indices to construct a phase space that enables robust discrimination between the topologies of molecules of different types and sizes. The necessity of our approach is illustrated with two practical examples. First, we use the phase space to distinguish between topologies of RNA secondary structures with different degrees of branching and to show how different tree representations of RNA influence the resulting topology. Second, we compare two methods of coarse-graining the structures of branched macromolecules and use the phase space of normalized topological indices to highlight the differences between the two methods. Taken together, our work establishes a phase space based on normalized topological indices that allows for a well-defined way of discriminating between branching architectures of different macromolecules as well as of methods of coarse-graining them.
@@
II Phase space of normalized topological indices
II.1 Tree representation of branched macromolecules
Branching is characteristic of macromolecules across different systems and scales (Fig. 1). To compare their topologies, one can map their structure onto a mathematical tree $T$ with $N$ edges (the tree size), with $V(T)$ denoting the set of its nodes and $E(T)$ denoting the set of its edges. This mapping is generally not unique and depends on the macromolecule [44]. For instance, a molecular (chemical) graph has atoms as nodes and edges that represent chemical bonds (Fig. 1a). Due to the nature of carbon bonding, chemical graphs have a maximum branching degree of $4$ regardless of their size. For polymer structures, a typical tree representation is one where a node is a collection of atoms at a polymer branch point and an edge is a linear structure formed by a chain of bonds between two such branch points (Fig. 1b). A more complex case, RNA is a linear polymer whose constituent nucleotides interact and base-pair with one another and give rise to a complex and highly branched secondary structure. This structure is often mapped to a tree by associating the base-paired (double-stranded) regions with edges and unpaired (single-stranded) regions with nodes (Fig. 1c). As we will see later on, such a coarse-grained fashion of mapping macromolecules to trees can lead to RNA trees with significantly larger branching degrees compared to chemical graphs.
II.2 Constructing the phase space of topological indices
Once a branched macromolecule is mapped to a mathematical tree, its architecture can be studied using methods from graph theory. This includes a wide variety of topological indices designed to describe the topology of trees and, more generally, graphs, with applications ranging from the description of chemical properties of molecules to connectivity of networks [45, 46, 47]. Broadly speaking, we can divide topological indices into path-, degree-, centrality-, spectrum-, and information-based indices depending on the main graph property used in their definition. For simplicity, we focus here on $18$ commonly used topological indices; their definitions are given in Supporting Information (SI). Since most topological indices were designed to be applied to graphs in general—which include cycles—they tend to be highly correlated when used on trees (Fig. S1 in SI). Individual indices are also highly degenerate and do not discriminate well between different tree topologies [51]. Therefore, our first aim is to select a pair of topological indices which are correlated as little as possible so that we can use their phase space for comparison of tree topologies.
We find that out of the $18$ topological indices we examined, one of the least correlated pairs (Fig. S1) is formed by the second-order network coherence $H_{2}$ and the atom bond connectivity index $\mathrm{ABC}$. The second-order network coherence is spectrum-based and defined as
@@ II.2 Constructing the phase space of topological indices
The phase space formed by the pair of indices $(\mathrm{ABC},H_{2})$ permits us to discriminate between tree topologies at fixed tree size $N$ (Fig. 2a) significantly better than if we would have used either of the two indices alone or a different combination of two indices such as the commonly used Wiener and Randić indices $W$ and $R$ (Fig. S6 in SI). We can furthermore trace the boundary of the $(\mathrm{ABC},H_{2})$ phase space with special cases of tree topologies, ranging from linear and dumbbell trees to star trees, star-like trees, and spider trees (Fig. 2b), which provides us with a basic understanding of topologies found in different regions of the phase space.
Figure 2: Phase space constructed from topological indices ABC and $H_{2}$. (a) $(\mathrm{ABC},H_{2})$ space of $10^{4}$ random trees with $N=16$ edges. Distributions at the axes’ boundaries show the probability distribution functions of individual indices. (b) Different tree topologies that trace the boundary of the phase space, with large symbols in panels (a) and (d) corresponding to the shown tree structures. The trees at the boundary range from a linear tree, dumbbell trees, star tree, star-like trees, spider tree, and back to a linear tree; their construction is shown in Fig. S5 in SI. (c) $(\mathrm{ABC},H_{2})$ space for trees of four different sizes, where $10^{4}$ random trees are sampled for each $N$. (d) $(\mathrm{ABC}^{\prime},H_{2}^{\prime})$ space of trivially normalized indices [Eq. (3)] showing trees of four different sizes with $10^{4}$ random trees for each size (cf. panel (c)). (e) $(\mathrm{ABC}^{*},H_{2}^{*})$ space of indices normalized through their cumulative distribution functions [Eq. (5)] for $10^{4}$ random trees with $N=16$. Distributions at the axes’ boundaries show the probability distribution functions of individual indices. (f) Heatmaps of the probability density of trees in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space for $10^{4}$ random trees of size $N=30$, $50$, $100$, and $200$. (g) $(\mathrm{ABC}^{*},H_{2}^{*})$ space of all non-isomorphic trees with $N=16$, showing separately chemical and caterpillar trees as well as their union.
II.3 Comparing topologies of trees of different sizes
@@
Such normalization limits an index to the range $\mathrm{TI}^{\prime}\in[0,1]$, where $\mathrm{TI}^{\prime}(\mathrm{linear})=0$ and $\mathrm{TI}^{\prime}(\mathrm{star})=1$. However, this turns out to be too simple for the purpose of comparison, as it progressively compresses the phase space as $N$ increases (Fig. 2d) and makes discrimination between different branching architectures difficult. The compression of the phase space is due to the mean values of a topological index and its values for linear and star trees all scaling differently with tree size (Table S1 and Fig. S2). In our case, this effect is much more pronounced for $H_{2}$ compared to $\mathrm{ABC}$.
@@
To determine the $\operatorname{cdf}(\mathrm{TI})$ of a given topological index, we need to first obtain its probability density function, $\operatorname{pdf}(\mathrm{TI})$. This is not a trivial endeavour since probability density functions for topological indices are not known. We therefore tested different continuous parametric distributions and chose the best-fitting one for each topological index (see Appendix and Fig. S3 in SI). The $\operatorname{pdf}(\mathrm{ABC})$ is best described by a log-normal distribution [Eq. (6)], while the Moyal distribution [Eq. (7)] is the best choice for the $\operatorname{pdf}(H_{2})$. The $N$-dependence of the fit parameters for both distributions is well-described by a simple power law and allows us to use the fitted $\operatorname{pdf}(\mathrm{TI})$ to immediately determine the cumulative distribution function $\operatorname{cdf}(\mathrm{TI})$ for an arbitrary tree size $N$.
The tree topologies assuming the extreme values of $\operatorname{cdf}(\mathrm{TI}(T))$ correspond to the tree topologies that assume the extreme values of the original index $\mathrm{TI}$. If we want to define a normalized index $\mathrm{TI}^{*}$ so that its values for the linear and the star tree are fixed to be $\mathrm{TI}^{*}(\mathrm{linear})=0$ and $\mathrm{TI}^{*}(\mathrm{star})=1$, we have to normalize the cumulative distribution function accordingly. We thus define a normalized topological index $\mathrm{TI}^{*}$ as:
@@
II.4 Properties of $(\mathrm{ABC}^{*},H_{2}^{*})$ space
We can now use the normalization via cumulative distribution functions to create a space of normalized indices $(\mathrm{ABC}^{*},H_{2}^{*})$ which leads to a significantly more homogeneous phase space and enables us to compare topologies of trees of arbitrary size (panels (e) and (f) of Fig. 2). The $(\mathrm{ABC}^{*},H_{2}^{*})$ space also allows us to interpret the relationships between different tree topologies at its boundary (Fig. 2e) as well as inside it. For instance, chemical trees, which have a maximum branching degree of $4$, and caterpillar trees, whose nodes all lie within distance $1$ of a central path, occupy distinct parts of the phase space (Fig. 2g). The part of the phase space that chemical and caterpillar trees occupy furthermore shrinks with increasing tree size (panels (a) and (b) of Fig. 3) since their overall proportion compared to all trees decreases with increasing $N$. More generally, the position and proportion of different types of trees in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space is broadly determined by their maximum node degree and number of degree $2$ nodes. Specifically, the maximum node degree of a tree in the phase space increases as we go from a linear tree at $(0,0)$ to a star tree at $(1,1)$ (Fig. 3c); at the same time, the number of degree $2$ nodes diminishes (Fig. 3d).
Figure 3: Characterization of the $(\textrm{ABC}^{*},H_{2}^{*})$ space. Changes in the proportion of (a) chemical and (b) caterpillar trees in the phase space with increasing tree size $N$, estimated from samples of $10^{4}$ random trees. Heatmaps of (c) maximum node degree and (d) fraction of degree $2$ nodes for trees with $N=16$. (e)(h): Changes in the phase space for trees with $N=50$ upon performing four different types of tree transformations $10^{3}$ times on five randomly chosen trees (black circles): random leaf move, random leaf remove, Prüfer shuffle, and random subtree swap. See the main text and Appendix for details on these transformations.
We can also use the pair of normalized indices $(\mathrm{ABC}^{*},H_{2}^{*})$ to examine how different transformations of trees reposition their topology in the phase space. If we either move or remove a random leaf node, this repositions the new tree topology along the $\mathrm{ABC}$ axis but leaves its position along the $H_{2}$ axis fairly unchanged (panels (e) and (f) of Fig. 3). Transformations that affect larger parts of the tree topology—such as the Prüfer shuffle, which retains the node degree distribution of a tree but rearranges the connections between nodes, and random subtree swap, where an edge is randomly deleted and a new one randomly added—change the position of tree topology in the phase space more drastically (panels (g) and (h) of Fig. 3).
@@ II.4 Properties of $(\mathrm{ABC}^{*},H_{2}^{*})$ space
III Example: different tree representations of RNA structure
Base-pairing between the four constituent nucleotides of RNA leads to a branched secondary structure [21, 22] with numerous high-degree branching nodes ($d(i)>4$[20], making RNA unique among known branched macromolecules. This often carries functional significance: For instance, highly-branched RNA genomes of some viruses are compact and consequently more easily encapsidated during virion formation than less branched ones [56, 57]. Graphs present a natural way to coarse-grain RNA structure and analyze its branching topology [58]. However, the choice of the mapping of the base-pair pattern onto a mathematical tree is not straightforward and several options exist, designed with different purposes in mind [59, 60]. Perhaps the simplest way to map RNA secondary structure to a tree is by mapping double-stranded regions (base pairs) to edges, while single-stranded regions (unpaired nucleotides) are mapped to nodes connecting the edges [21, 22]. An alternative way, used in the RNA-as-graph (RAG) framework [58], omits from nodes those unpaired regions which consist of a single unpaired nucleotide or a single non-complementary pair of nucleotides, and does not consider lonely base pairs as edges. (For details on the two approaches, see Appendix.)
@@ III Example: different tree representations of RNA structure
Secondary structures of the three RNA sequences are sampled from their thermal ensembles (Methods and Appendix) and then mapped to trees using either the simple or the RAG method. While the RAG method (Fig. 4d) results in trees on average two thirds the size of those obtained using the simple method (Fig. 4a), the $(\mathrm{ABC}^{*},H_{2}^{*})$ space allows us to compare the RNA topologies resulting from the two mapping methods (panels (b) and (e) of Fig. 4). For both methods we observe that the three RNA sequences are separated in the phase space along the $H_{2}^{*}$ axis. The compact BlShV genome occupies large $H_{2}^{*}$ values, the non-compact BVQ genome occupies small $H_{2}^{*}$ values, and the random RNA sequence is located in-between the two. The $\mathrm{ABC}^{*}$ index, on the other hand, shows the differences between the two mapping methods. For the simple method, most trees acquire values of $\mathrm{ABC}^{*}$ close to zero, showing little fluctuation between the structures sampled from the thermal ensemble. For the RAG method, however, the $\mathrm{ABC}^{*}$ values of both RNA genomes are distributed along the entire range of values, implying larger thermal fluctuations between the structures; $\mathrm{ABC}^{*}$ values of the random RNA, while non-zero, remain small and in this way differ from both biological RNAs. These differences can be traced to the proportion of degree $2$ nodes produced by the two mapping methods, as the simple method (Fig. 4c) results in approximately twice as many degree $2$ nodes than the RAG method (Fig. 4f). As we have already observed previously, the amount of degree $2$ nodes is an important factor that determines the position of a tree topology in the phase space (Fig. 3).
@@
IV Example: Coarse-graining the topologies of random trees